OpenStudy (anonymous):

If ƒ(x ) = 3x + 1, then ƒ(a + h ) - ƒ(a ) = h 3h 3h + 2

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

@magepker728

OpenStudy (anonymous):

i think ur missing something

OpenStudy (anonymous):

I'm not

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

3h

OpenStudy (anonymous):

\[If f(x) = 3x+1 and f ^{-1}=\frac{ x-1 }{ 3 } then f(2)=\]

OpenStudy (anonymous):

@joyraheb I couldn't figure out how to put this into a new post

OpenStudy (anonymous):

???

OpenStudy (cggurumanjunath):